# ProphetKV: User-Query-Driven Selective Recomputation for Efficient KV Cache Reuse in Retrieval-Augmented Generation

### v1

@@ E.3 Accuracy Results on other LongBench datasets
For challenging cases such as 2wikimqa (WQA) and hotpotqa (HQA), Naive Reuse exhibits a significant accuracy drop relative to full recomputation. This suggests that these tasks require a more comprehensive understanding of context and cross-chunk information interaction. In these cases, our ProphetKV achieves results comparable to full recomputation, demonstrating its effectiveness in preserving critical information for accurate response generation. Conversely, for simpler cases such as gov_report (GRep), Naive Reuse achieves performance comparable to full recomputation, indicating that these tasks may depend more on local context and less on cross-chunk interactions. In such scenarios, all partial recomputation methods, including ProphetKV, perform well, highlighting their ability to maintain accuracy while reducing computational overhead.
We report results only for Llama-3.1-8B-Inst. and Qwen2.5-14B-Inst. due to time constraints. Qwen3-14B Thk. requires a significantly longer time to generate answers, as the input length approaches 16k tokens, and the generation length is set to 4k tokens to ensure answer completeness.

### v2

@@ E.3 Accuracy Results on other LongBench datasets
Qwen-3-14B Thk. 75.52 67.82 8.87 17.93 29.05 7.12 22.14 50.44 65.79 33.33 99.75 42.23 9.50 47.40 14.33 39.41
NaiveReuse 32.12 22.67 7.08 5.94 17.81 1.63 21.24 30.83 31.52 0.81 13.75 32.82 1.72 36.05 1.61 17.17
CacheBlend 65.76 59.45 8.41 15.39 29.45 4.52 22.12 42.31 58.14 3.25 66.75 40.83 3.47 47.42 2.53 31.32
EPIC 63.80 56.84 8.93 15.01 29.34 3.25 22.21 45.17 59.10 3.25 57.83 40.42 3.48 47.20 2.24 30.54
KVShare 68.53 60.69 8.26 15.28 29.31 3.75 22.02 43.12 57.05 2.44 57.50 41.85 2.99 48.18 2.60 30.90
ProphetKV 72.22 67.74 8.51 15.52 29.03 6.45 21.94 44.54 56.27 3.25 99.75 39.19 4.00 49.31 4.47 34.81
For challenging cases such as 2wikimqa (WQA) and hotpotqa (HQA), Naive Reuse exhibits a significant accuracy degradation relative to full recomputation. This observation suggests that these tasks require a more comprehensive understanding of global context and cross-chunk interactions. In contrast, ProphetKV achieves performance comparable to full recomputation in these settings, demonstrating its effectiveness in preserving critical information necessary for accurate response generation. Conversely, for simpler cases such as gov_report (GRep), Naive Reuse attains performance on par with full recomputation, indicating that these tasks primarily rely on local context and are less sensitive to cross-chunk interactions. In such scenarios, all partial recomputation methods, including ProphetKV, perform well, highlighting their ability to maintain accuracy while reducing computational overhead. Finally, certain cases in Qwen-3-14B Thk. (e.g., passage_count (PassCnt) and repobench-p (RepoB)) suffer from excessively long thinking generation lengths, causing the model to exceed the maximum length of 4K tokens during answer generation. Consequently, the scores for these cases are substantially lower than those of other datasets.
@@
Qwen-3-14B Thk. 8.32 29.59 9.07 20.86 50.44 32.09 38.55 9.30 24.78
NaiveReuse 7.03 21.89 1.27 16.33 30.83 3.92 31.15 1.95 14.30
CacheBlend 8.44 29.49 2.55 20.18 42.31 11.15 36.78 3.16 19.26
EPIC 8.29 29.52 2.61 20.22 45.17 8.78 36.90 2.08 19.20
KVShare 8.26 29.59 1.98 20.17 43.12 7.43 37.96 3.11 18.95
ProphetKV 8.62 29.42 6.27 20.24 44.54 6.08 37.02 3.20 19.42